OpenStudy (anonymous):

Can you please help me with this question? Solve the logarithmic equation algebraically. Approximate the result to three decimal places. (If there is no solution, enter NO SOLUTION.) ln(x + 5) – ln(x – 3) = ln(x)

OpenStudy (anonymous):

I think you meant log a - log b = log(a/b)

OpenStudy (anonymous):

But when you subtract, isn't the original one supposed to be in division form? log(a/b) = log (a) - log (b)

OpenStudy (anonymous):

@tcarroll010 I used the law but it didnt work

OpenStudy (saifoo.khan):

OMG! *KILL ME*

OpenStudy (saifoo.khan):

thanks for the correction. @tcarroll010

OpenStudy (anonymous):

no sweat, all you did was a mental slip. I did one earlier. We all do.

OpenStudy (saifoo.khan):

(:

OpenStudy (anonymous):

And then, ln [(x + 5)/(x - 3)] = ln 3 Then put e^ for each side. e^(ln x) = x

OpenStudy (anonymous):

Can you please show me the steps?

OpenStudy (anonymous):

e^(ln [(x + 5)/(x - 3)]) = e^(ln 3) gives (x + 5)/(x - 3) = 3 and then you can maybe solve from there.

OpenStudy (anonymous):

I made a typo: ln [(x + 5)/(x - 3)] = ln x Then put e^ for each side. e^(ln [(x + 5)/(x - 3)]) = e^(ln x) gives (x + 5)/(x - 3) = x and then you can maybe solve from there.

OpenStudy (anonymous):

i did, i got x=5 and -1

OpenStudy (anonymous):

Good job! Well done! Even with my initial typo!

OpenStudy (anonymous):

but I checked they were not right though

OpenStudy (anonymous):

That's right, you can only have x = 5 because you can't have a ln of a negative number.

OpenStudy (anonymous):

oh ok, thank you! will put that in and see what happens!

OpenStudy (anonymous):

When we came down to the equation (x + 5)/(x - 3) = x we came to that by power of e. The original equation had a ln x, so we have to put a restriction on x to be positive.

OpenStudy (anonymous):

thank god, thank you so much carrol! I got it right, I just had a chance left, if I put in somethign instead of 5 then there would've be no more chances left for me to try other solutions!

OpenStudy (anonymous):

Well that worked out well for you!

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

And you are quite welcome!